OpenStudy (anonymous):

The velocity of a particle moving along the x-axis is given, for t>0, by v=(50.0t - 2.0t^3) m/s, where t is in seconds. What is the acceleration of the particle when (after t=0) it achieves its maximum displacement in the positive x direction?

OpenStudy (amistre64):

since velocity is the derivative of displacement, when v = 0, we have critical points to test for displacement. the derivative of velocity is acceleration, so when you determine the value of t, plug it in the acceleration function

OpenStudy (amistre64):

\[v(t) = 50t - 2t^3\] \[2t(25 - t^2)=0\] so apparently when t= {-5,0,5}; which is only valid for t=5 \[v'(t) = a(t) =50 - 6t^2\]

OpenStudy (amistre64):

something feels wierd there, but i cant see an error in my thought

OpenStudy (amistre64):

if velocity is zero, acceleration would be zero .... is what im usually thinking :/

OpenStudy (anonymous):

Thank you for you help! I haven't been able to check your work against mine yet...but i do appreciate your effort!

OpenStudy (amistre64):

youre welcome. i do wish i was more confident in the results :) Im pretty sure it is because we started with a cubic for velocity, or im just missing something.